OpenStudy (anonymous):

Combine like terms: (7y2 - 3y) +8y2 - 8y)

OpenStudy (anonymous):

The like terms are: 7y^2 and 8y^2 -3y and -8y. Add them. 7y^2 + 8y^2 = 15y^2 -3x + (-8y) = -3y - 8y = -11y So the answer would be 15y^2 - 11y .

OpenStudy (anonymous):

i got that answer so perfect!